OpenStudy (anonymous):

Find a number that is between 6/11 and 0.67(0.67 repeats) . A. 11/18 B. 69/99 C. 24/33 D. 26/55

Parth (parthkohli):

Add both and divide by 2.

OpenStudy (anonymous):

To find 6/11 as a fraction, just divide 6 by 11, which is 0.54 repeating

Parth (parthkohli):

Or represent 0.676767... as a fraction.

OpenStudy (anonymous):

To see which numbers are inbetween, you'll have to do the same to the answer choices

Parth (parthkohli):

6/11 and 67/99. 66/99 and 67/99.

Parth (parthkohli):

Now it's all checking.

OpenStudy (anonymous):

A. A. 11/18= 0.61 B. 69/99= 0.69 repeating C. 24/33= 0.72 repeating D. 26/55= 472727272 so on (I think this is the easy way)

OpenStudy (anonymous):

All you have to do now is see what numbers are inbetween 0.54 and 0.67 :)

Parth (parthkohli):

Oh wow, I messed up big time.

OpenStudy (anonymous):

So, what do you think the answer is ?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Right :)